A galvanometer has a full scale defection of 1.0A and internal resistance of 0.5ohms, if it is connected to a current of 30 A, calculate the value of resistance of shunt required with the meter without damaging it
1
Expert's answer
2019-08-23T09:34:59-0400

The maximum voltage in the galvanometer resistor is given by:


"V_{G}=R_{in}*I_{G}"


Where;

  • Electric current "I_{G}=1.0A"
  • electric resistance "R_{in}=0.5\Omega"


Assessing numerically.


"V_{G}=1.0A*0.5\Omega=0.5V"


Parallel resistance has the same potential difference.


"V_{G}=V_{shunt}=0.5V"


The current flowing must be equal to:"I_{shunt}=30A-1.0A=29A"


Therefore the value of R is calculated with Ohm's Law


"R_{shunt}=\frac{V_{shunt}}{I_{shunt}}"


Numerically evaluating


"R_{shunt}=\frac{0.5V}{29A}=1.72*10^{-2}\Omega"


the value of resistance of shunt required with the meter without damaging it is


"\boxed{R_{shunt}=1.72*10^{-2}\Omega}"